MAT 205 CALCULUS III Vector Calculus SPRING 2000

This course represents the third semester in the standard Calculus sequence. We will develop the theory and practice of differentiation and integration for functions of several variables, building upon your previous experience with functions of one variable. There will be a balance of theory and practice, although formal proofs will not be emphasized.

PREREQUISITES: Either MAT 127, or MAT 132, or MAT 142 or level 9 at the mathematics placement examination.

LECTURES: Tu-Th 2:20–3:40, Physics P 117

INSTRUCTOR: Mark de Cataldo; Office: MAT 3-115 (math building, third floor, room 115); Phone: 2-8262; e-mail: mde@math.sunysb.edu Office hours: Tu 5-6, Th 1-2, or by appointment.

TEXTBOOK: Multivariable Calculus: concepts and contexts, by James Stewart, Brooks/Cole, 1998.

RECITATIONS: Fr 12:40-1:35, ROOM SB Union 214

TEACHING ASSISTANT: Jason Van Hulse; Office: MAT 3-104 (math building, third floor, room 104); e-mail: hultz@math.sunysb.edu Office hours: To Be Announced

TENTATIVE SCHEDULE OF EXAMS

MIDTERM 1: Tuesday February 29, 2000; 2:20 – 3:40 pm, in class.

MIDTERM 2: Tuesday April 4, 2000; 2:20 – 3:40 pm, in class.

FINAL: Thursday, May 11, 2000, NOON-3 pm; unless otherwise announced, the examination will be given in the room where the regular class is held.

IMPORTANT. Make sure that you will be available at these times. There will be no make-ups for missed mid-terms. However, if you need to miss an exam for an acceptable reason, and provide me with an acceptable written excuse, the relevant mid-term will be ‘dropped’ (ignored) in computing your course grade. Incompletes will be granted only if documented circumstances beyond your control prevent you from taking the final exam.

HOMEWORK: Usually assigned in class on Tuesdays, and usually due the following Tuesday. Only some of the problems will be graded, but which ones will not be announced in advance. You are therefore strongly advised to do all the problems. Your ten best homework scores will be counted toward your course grade. Please note that late homework will not be accepted.

QUIZZES: Quizzes will be given twice a month during recitation.

GRADE: Midterm 1 = 20%, Midterm 2 = 20%, Final= 40%, Homework = 10%, Quizzes = 10%.

CONTACTING THE INSTRUCTOR AND/OR THE TEACHING ASSISTANT. E-mail is the best method for making appointments outside normal office hours. E-mail is not, however, a good way to ask math questions, as our typing abilities are very limited.

SPECIAL NEEDS. If you have a physical, psychiatric, medical, or learning disability that may affect your ability to carry out the assigned course work, please contact the office of Disabled Student Services (DSS), Humanities Building, room 133, telephone 632-6748/TDD. DSS will review your concerns and determine what accomodations may be necessary and appropriate. All information and documentation of disability is confidential.